### v1

# Wasserstein Transfer Learning

Abstract
Transfer learning is a powerful paradigm for leveraging knowledge from source domains to enhance learning in a target domain. However, traditional transfer learning approaches often focus on scalar or multivariate data within Euclidean spaces, limiting their applicability to complex data structures such as probability distributions. To address this, we introduce a novel framework for transfer learning in regression models, where outputs are probability distributions residing in the Wasserstein space. When the informative subset of transferable source domains is known, we propose an estimator with provable asymptotic convergence rates, quantifying the impact of domain similarity on transfer efficiency. For cases where the informative subset is unknown, we develop a data-driven transfer learning procedure designed to mitigate negative transfer. The proposed methods are supported by rigorous theoretical analysis and are validated through extensive simulations and real-world applications.
Keywords: Data Aggregation, Domain Adaptation, Multitask Learning, non-Euclidean Output, Optimal Transport
1 Introduction
In recent years, transfer learning (Torrey & Shavlik 2010) has emerged as a powerful paradigm in machine learning, enabling models to leverage knowledge acquired from one domain and apply it to related tasks in another. This approach has proven especially valuable in scenarios where data collection and labeling can be costly, or where tasks exhibit inherent similarities in structure or representation. While early successes focused on conventional data types such as images (Shin et al. 2016), text (Raffel et al. 2020), and tabular data (Weiss et al. 2016), there is growing interest in extending these methods to more complex data structures. Such data often reside in non-Euclidean spaces and lack basic algebraic operations like addition, subtraction, or scalar multiplication, posing challenges for traditional learning algorithms. A key example is probability distributions (Petersen et al. 2022), where for example the sum of two density functions does not yield a valid density.
Samples of univariate probability distributions are increasingly encountered across various research domains, such as mortality analysis (Ghodrati & Panaretos 2022), temperature studies (Zhu & Müller 2023), and physical activity monitoring (Lin et al. 2023), among others (Petersen et al. 2022). Recently, there has been a growing focus on directly modeling distributions as elements of the Wasserstein space, a geodesic metric space related to optimal transport (Villani 2009, Panaretos & Zemel 2020). The absence of a linear structure in this space motivates the development of specialized transfer learning techniques that respect its intrinsic geometry.
To address this gap, we introduce Wasserstein Transfer Learning (WaTL), a novel transfer learning framework for regression models where outputs are univariate probability distributions. WaTL effectively leverages knowledge from source domains to improve learning in a target domain by intrinsically incorporating the Wasserstein metric, which provides a natural way to measure discrepancies between probability distributions.
1.1 Contributions
The primary contributions of this work are summarized as follows:
Methodology. We propose a novel transfer learning framework for regression models with distributional outputs, addressing the challenges inherent in the Wasserstein space, which lacks a conventional linear structure. Our framework includes an efficient algorithm for cases where the informative subset of source domains is known, and a data-driven algorithm for scenarios where the subset is unknown. To the best of our knowledge, this is the first comprehensive transfer learning approach specifically designed for regression models with outputs residing in the Wasserstein space.
Theoretical analysis. We establish the asymptotic convergence rates for the WaTL algorithm in both the case where the informative set is known and the more challenging scenario where it must be estimated. In the latter case, we also prove that the informative set can be consistently identified. In both settings, we demonstrate that WaTL effectively improves model performance on the target data by leveraging information from the source domain. The proofs rely heavily on empirical process theory and a careful analysis of the covariate structure. Our key theoretical results extend beyond responses lying in the Wasserstein space, offering potential applications to other complex outputs.
Simulation studies and real-world applications. We evaluate WaTL through simulations and real data applications, demonstrating its effectiveness in improving target model performance by leveraging source domain information. The benefits become more pronounced with larger source sample sizes, underscoring its ability to harness transferable knowledge.
1.2 Related Work
Transfer learning. Transfer learning aims to improve performance in a target population by leveraging information from a related source population and has seen wide application across domains (e.g., Huang et al. 2013, Gong et al. 2012, Choi et al. 2017, Zhou et al. 2024). Recent theoretical developments have focused on regression in Euclidean settings, including high-dimensional linear (Li et al. 2022) and generalized linear models (Tian & Feng 2023), nonparametric regression (Cai & Pu 2024, Lin & Reimherr 2024), and function mean estimation from discretely sampled data (Cai et al. 2024). In parallel, optimal transport has been used to measure distributional shifts for domain adaptation (Courty et al. 2017, Redko et al. 2017). However, to the best of our knowledge, no existing work has investigated transfer learning in regression models where outputs are probability distributions residing in the Wasserstein space. This represents a significant gap in the literature, highlighting the need for novel methodologies that address this challenging yet important setting.
Distributional data analysis. The increasing prevalence of data where distributions serve as fundamental units of observation has spurred the development of distributional data analysis (Petersen et al. 2022). Recent advancements in this field include geodesic principal component analysis in the Wasserstein space (Bigot et al. 2017), autoregressive models for time series of distributions (Zhang et al. 2022, Zhu & Müller 2023), and distribution-on-distribution regression (Ghodrati & Panaretos 2022, Chen et al. 2023). Leveraging the Wasserstein metric, regression models with distributional outputs and Euclidean inputs can be viewed as a special case of Fréchet regression (Petersen & Müller 2019), which extends linear and local linear regression to outputs residing in general metric spaces. In practical scenarios where only finite samples from the unknown distributional output are available, empirical measures have been utilized as substitutes for the unobservable distributions in regression models (Zhou & Müller 2024).
2 Preliminaries
2.1 Notations
[…]
We write $a_{n}\lesssim b_{n}$ if there exists a positive constant $C$ such that $a_{n}\leq Cb_{n}$ when $n$ is large enough and $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$. The notation $a_{n}=O_{p}(b_{n})$ implies that $P(|a_{n}/b_{n}|\leq C)\to 1$ for some constant $C>0$, while $a_{n}=o_{p}(b_{n})$ implies that $P(|a_{n}/b_{n}|>c)\to 0$ for any constant $c>0$. Superscripts typically indicate different data sources, while subscripts distinguish individual samples from the same source.
2.2 Wasserstein Space
[…]
The space $\mathcal{W}$, endowed with the Wasserstein metric, forms a complete and separable metric space, commonly known as the Wasserstein space (Villani 2009).
Assuming $E\{d_{\mathcal{W}}^{2}(\nu,\mu)\}<\infty$ for all $\mu\in\mathcal{W}$, the Fréchet mean (Fréchet 1948) of a random distribution $\nu\in\mathcal{W}$ is given by $\nu_{\oplus}=\operatorname*{arg\,min}_{\mu\in\mathcal{W}}E\{d_{\mathcal{W}}^{2}(\nu,\mu)\}$. Since the Wasserstein space $\mathcal{W}$ is a Hadamard space (Kloeckner 2010), the Fréchet mean is well-defined and unique. Moreover, from (1), it follows that the quantile function of the Fréchet mean, denoted as $F^{-1}_{\nu_{\oplus}}$, satisfies $F^{-1}_{\nu_{\oplus}}(u)=E\{F_{\nu}^{-1}(u)\},u\in(0,1)$.
2.3 Fréchet Regression
Consider a random pair $(X,\nu)$ with joint distribution $\mathcal{F}$ on the product space $\mathbb{R}^{p}\times\mathcal{W}$. Let $X$ have mean $\theta=E(X)$ and covariance matrix $\Sigma=\operatorname{Var}(X)$, where $\Sigma$ is assumed to be positive definite. To establish a regression framework for predicting the distributional response $\nu$ from the covariate $X$, we employ the Fréchet regression model, which extends multiple linear regression and local linear regression to scenarios where responses reside in a metric space (Petersen & Müller 2019). The Fréchet regression function is defined as the conditional Fréchet mean of $\nu$ given $X=x$,
$m(x)=\operatorname*{arg\,min}_{\mu\in\mathcal{W}}E\{d_{\mathcal{W}}^{2}(\nu,\mu)|X=x\}.$
For a detailed exposition of Fréchet regression, we refer the reader to Petersen & Müller (2019). Given $n$ independent realizations $\{(X_{i},\nu_{i})\}_{i=1}^{n}$, we define the empirical mean and covariance of $X$ as $\overline{X}=\frac{1}{n}\sum_{i=1}^{n}X_{i}$ and $\widehat{\Sigma}=\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\overline{X})(X_{i}-\overline{X})^{\mathrm{\scriptscriptstyle T}}$.
The global Fréchet regression extends classical multiple linear regression and estimates the conditional Fréchet mean as
$m_{G}(x)=\operatorname*{arg\,min}_{\mu\in\mathcal{W}}E\{s_{G}(x)d_{\mathcal{W}}^{2}(\nu,\mu)\},$
where the weight function is given by $s_{G}(x)=1+(X-\theta)^{\mathrm{\scriptscriptstyle T}}\Sigma^{-1}(x-\theta)$. The empirical estimator is formulated as
$\widehat{m}_{G}(x)=\operatorname*{arg\,min}_{\mu\in\mathcal{W}}\frac{1}{n}\sum_{i=1}^{n}s_{iG}(x)d_{\mathcal{W}}^{2}(\nu_{i},\mu),$
where $s_{iG}(x)=1+(X_{i}-\overline{X})^{\mathrm{\scriptscriptstyle T}}\widehat{\Sigma}^{-1}(x-\overline{X})$.
Similarly, local Fréchet regression extends classical local linear regression to settings with metric space-valued outputs. In the case of a scalar predictor $X\in\mathbb{R}$, the local Fréchet regression function is
$m_{L,h}(x)=\operatorname*{arg\,min}_{\mu\in\mathcal{W}}E\{s_{L}(x,h)d_{\mathcal{W}}^{2}(\nu,\mu)\},$
where the weight function is $s_{L}(x,h)=K_{h}(X-x)\{u_{2}-u_{1}(X-x)\}/\sigma_{0}^{2}$, with $u_{j}=E\{K_{h}(X-x)(X-x)^{j}\},j=0,1,2$, and $\sigma_{0}^{2}=u_{0}u_{2}-u_{1}^{2}$. Here, $K_{h}(\cdot)=h^{-1}K(\cdot/h)$ is a kernel function with bandwidth $h$. The empirical version is given by
$\widehat{m}_{L,h}(x)=\operatorname*{arg\,min}_{\mu\in\mathcal{W}}\frac{1}{n}\sum_{i=1}^{n}s_{iL}(x,h)d_{\mathcal{W}}^{2}(\nu_{i},\mu),$
where $s_{iL}(x,h)=K_{h}(X_{i}-x)\{\widehat{u}_{2}-\widehat{u}_{1}(X_{i}-x)\}/\widehat{\sigma}_{0}^{2}$, with $\widehat{u}_{j}=n^{-1}\sum_{i=1}^{n}K_{h}(X_{i}-x)(X_{i}-x)^{j},j=0,1,2$, and $\widehat{\sigma}_{0}^{2}=\widehat{u}_{0}\widehat{u}_{2}-\widehat{u}_{1}^{2}$.
3 Methodology
3.1 Setup
We consider a transfer learning problem where target data $\{(X_{i}^{(0)},\nu_{i}^{(0)})\}_{i=1}^{n_{0}}$ are sampled independently from the target population $(X^{(0)},\nu^{(0)})\sim\mathcal{F}_{0}$, and source data $\{(X_{i}^{(k)},\nu_{i}^{(k)})\}_{i=1}^{n_{k}}$ are sampled independently from the source population $(X^{(k)},\nu^{(k)})\sim\mathcal{F}_{k}$, for $k=1,\ldots,K$. The goal is to estimate the target model using both the target data and the source data from $K$ related studies.
For $k=0,\ldots,K$, assume $X^{(k)}$ has mean $\theta_{k}$ and covariance $\Sigma_{k}$, with $\Sigma_{k}$ positive definite. Define the empirical mean and covariance of $\{X_{i}^{(k)}\}_{i=1}^{n_{k}}$ as $\overline{X}_{k}=n_{k}^{-1}\sum_{i=1}^{n_{k}}X_{i}^{(k)}$ and $\widehat{\Sigma}_{k}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}(X_{i}^{(k)}-\overline{X}_{k})(X_{i}^{(k)}-\overline{X}_{k})^{\mathrm{\scriptscriptstyle T}}.$ For a fixed $x\in\mathbb{R}^{p}$, the weight function is $s_{G}^{(k)}(x)=1+(X^{(k)}-\theta_{k})^{\mathrm{\scriptscriptstyle T}}\Sigma_{k}^{-1}(x-\theta_{k}),$ with the sample version $s_{iG}^{(k)}(x)=1+(X_{i}^{(k)}-\overline{X}_{k})^{\mathrm{\scriptscriptstyle T}}\widehat{\Sigma}_{k}^{-1}(x-\overline{X}_{k}).$ The target regression function for a given $x\in\mathbb{R}^{p}$ is then $m_{G}^{(0)}(x)=\operatorname*{arg\,min}_{\mu\in\mathcal{W}}E\{s_{G}^{(0)}(x)d_{\mathcal{W}}^{2}(\mu,\nu^{(0)})\}.$ In the following, we present details on transfer learning for global Fréchet regression, where the key difference in the local Fréchet regression setting is the use of a different weight function. The technical details for transfer learning in local Fréchet regression are therefore deferred to Appendix D.
The set of informative auxiliary samples (informative set) consists of sources sufficiently similar to the target data. Formally, the informative set is defined as $\mathcal{A}_{\psi}=\big{\{}1\leq k\leq K:\|f^{(0)}(x)-f^{(k)}(x)\|_{2}\leq\psi\big{\}}$ for some $\psi>0$, where $f^{(k)}(x)=E\{s_{G}^{(k)}(x)F_{\nu^{(k)}}^{-1}\}$. For simplicity, let $n_{\mathcal{A}}=\sum_{k=1}^{K}n_{k}$.
3.2 Wasserstein Transfer Learning
We propose the Wasserstein Transfer Learning (WaTL) algorithm, which combines information from source datasets under the assumption that all source data are “informative enough.” This assumption implies that the discrepancies between the source and target are small enough to enhance estimation compared to using only the target. When this condition is met for all source datasets, the informative set is given by $\mathcal{A}_{\psi}=\{1,\ldots,K\}$. The detailed steps of WaTL are presented in Algorithm 1.
Algorithm 1 Wasserstein Transfer Learning (WaTL)
0:  Target and source data $\{(x_{i}^{(0)},\nu_{i}^{(0)})\}_{i=1}^{n_{0}}\cup\big{(}\cup_{1\leq k\leq K}\{(x_{i}^{(k)},\nu_{i}^{(k)})\}_{i=1}^{n_{k}}\big{)}$, regularization parameter $\lambda$, and query point $x\in\mathbb{R}^{p}$.
0:  Target estimator $\widehat{m}_{G}^{(0)}(x)$.
1:  Weighted auxiliary estimator: $\widehat{f}(x)=\frac{1}{n_{0}+n_{\mathcal{A}}}\sum_{k=0}^{K}n_{k}\widehat{f}^{(k)}(x)$, where $\widehat{f}^{(k)}(x)=n_{k}^{-1}\sum_{i=1}^{n_{k}}s_{iG}^{(k)}(x)F^{-1}_{\nu_{i}^{(k)}}$.
2:  Bias correction using target data: $\widehat{f}_{0}(x)=\operatorname*{arg\,min}_{g\in L^{2}(0,1)}\frac{1}{n_{0}}\sum_{i=1}^{n_{0}}s_{iG}^{(0)}(x)\|F^{-1}_{\nu_{i}^{(0)}}-g\|_{2}^{2}+\lambda\|g-\widehat{f}(x)\|_{2}.$
3:  Projection to Wasserstein space: $\widehat{m}^{(0)}_{G}(x)=\operatorname*{arg\,min}_{\mu\in\mathcal{W}}\big{\|}F^{-1}_{\mu}-\widehat{f}_{0}(x)\big{\|}_{2}.$
In Step 1, the initial estimate $\widehat{f}$ aggregates information from both the target and source, weighted by their respective sample numbers.
While this step incorporates valuable auxiliary information, the resulting estimate may be biased due to distributional differences between the target and source populations. In Step 2, the bias in $\widehat{f}$ is corrected by focusing on the target data. The regularization term $\lambda\|g-\widehat{f}(x)\|_{2}$ ensures a balance between target-specific precision and auxiliary-informed robustness. Theoretical guidelines for selecting $\lambda$ are provided in Theorem 2. The final step projects the corrected estimate $\widehat{f}_{0}$ onto the Wasserstein space, ensuring the output $\widehat{m}^{(0)}_{G}(x)$ satisfies the intrinsic geometry of $\mathcal{W}$. Such a projection exists and is unique as $\mathcal{W}$ is a closed and convex subset of $L^{2}(0,1)$.
3.3 Adaptive Selection of Informative Sources
In many practical scenarios, the assumption that all source datasets belong to the informative set $\mathcal{A}_{\psi}$ may not hold. To address this, we extend WaTL with an adaptive selection procedure to identify the informative set. The discrepancy for each source dataset $k$ is defined as $\psi_{k}=\|f^{(0)}(x)-f^{(k)}(x)\|_{2}$, which measures the distance between the target distribution and the auxiliary distribution. Since $f^{(0)}(x)$ and $f^{(k)}(x)$ are unknown, we compute an empirical estimate $\widehat{\psi}_{k}$ for $\psi_{k}$, which is used to adaptively estimate the informative set $\mathcal{A}_{\psi}$. To implement this approach, an additional input parameter $L$, which specifies the approximate number of informative sources, is required. In practice, $L$ can be treated as a tuning parameter and selected through cross-validation or other model selection techniques. The full procedure is formalized in Algorithm 2.
The proposed algorithm adaptively identifies the informative set $\widehat{\mathcal{A}}$ in Step 1 by evaluating the empirical discrepancy scores $\widehat{\psi}_{k}$. The selected set is then used to compute the weighted auxiliary estimator in Step 2, ensuring that only the most relevant source datasets contribute to the final target estimator. Steps 3 and 4 follow the same bias correction and projection procedures as described in Section 3.2. This adaptive approach enhances the robustness of WaTL by excluding irrelevant or highly dissimilar source datasets.
Algorithm 2 Adaptive Wasserstein Transfer Learning (AWaTL)
0:  Target and source data $\{(x_{i}^{(0)},\nu_{i}^{(0)})\}_{i=1}^{n_{0}}\cup\big{(}\cup_{1\leq k\leq K}\{(x_{i}^{(k)},\nu_{i}^{(k)})\}_{i=1}^{n_{k}}\big{)}$, regularization parameter $\lambda$, number of informative sources $L$, and query point $x\in\mathbb{R}^{p}$.
0:  Target estimator $\widehat{m}_{G}^{(0)}(x)$.
1:  Compute discrepancy scores. For each source dataset $k=1,\ldots,K$, compute the empirical discrepancy: $\widehat{\psi}_{k}=\|\widehat{f}^{(0)}(x)-\widehat{f}^{(k)}(x)\|_{2},$ where $\widehat{f}^{(k)}(x)=n_{k}^{-1}\sum_{i=1}^{n_{k}}s_{iG}^{(k)}(x)F^{-1}_{\nu_{i}^{(k)}}$. Construct the adaptive informative set by selecting the $L$ smallest discrepancy scores $\widehat{\mathcal{A}}=\{1\leq k\leq K:\widehat{\psi}_{k}\text{ is among the smallest }L\text{ values}\}.$
2:  Weighted auxiliary estimator: $\widehat{f}(x)=\frac{1}{\sum_{k\in\widehat{\mathcal{A}}\cup\{0\}}n_{k}}\sum_{k\in\widehat{\mathcal{A}}\cup\{0\}}n_{k}\widehat{f}^{(k)}(x).$
3:  Bias correction using target data: $\widehat{f}_{0}(x)=\operatorname*{arg\,min}_{g\in L^{2}(0,1)}\frac{1}{n_{0}}\sum_{i=1}^{n_{0}}s_{iG}^{(0)}(x)\|F^{-1}_{\nu_{i}^{(0)}}-g\|_{2}^{2}+\lambda\|g-\widehat{f}(x)\|_{2}.$
4:  Projection to Wasserstein space: $\widehat{m}_{G}^{(0)}(x)=\operatorname*{arg\,min}_{\mu\in\mathcal{W}}\big{\|}F^{-1}_{\mu}-\widehat{f}_{0}(x)\big{\|}_{2}.$
4 Theory
In this section, we establish the theoretical guarantees of the proposed WaTL and AWaTL algorithms using techniques from empirical process theory (van der Vaart & Wellner 2023). For WaTL, we present the following lemma, which characterizes the convergence rate for each term contributing to the weighted auxiliary estimator $\widehat{f}(x)$ computed in Step 1.
Condition 1.
For $k=0,\ldots,K$, the covariate $X^{(k)}$ is subguassian with $\|X^{(k)}\|_{\Psi_{2}}\in[\sigma_{1},\sigma_{2}]$, the mean vector satisfies $\|\theta_{k}\|\leq R_{1}$, and the spectrum of the covariance matrix $\Sigma_{k}$ lies within the interval $[R_{2},R_{3}]$. Moreover, $\nu^{(k)}$ is supported on a bounded interval.
Lemma 1.
Let $\widehat{f}^{(k)}(x)=n_{k}^{-1}\sum_{i=1}^{n_{k}}s_{iG}^{(k)}(x)F_{\nu_{i}^{(k)}}^{-1}$ and its population counterpart $f^{(k)}(x)=E\{s_{G}^{(k)}(x)F_{\nu^{(k)}}^{-1}\}$ for $k=0,\ldots,K$. Then under Condition 1, $\|\widehat{f}^{(k)}(x)-f^{(k)}(x)\|_{2}=O_{p}(n_{k}^{-1/2})$.
To derive the convergence rate of $\widehat{f}(x)$, we rely on the following condition.
Condition 2.
There exist positive constants $C_{1},C_{2},C_{3}$, $C_{4}$ such that
$\sum_{k=0}^{K}2e^{-C_{1}\frac{4}{R_{3}^{2}}n_{k}}+C_{2}e^{-C_{3}\big{(}\frac{2}{R_{3}}-\sqrt{\frac{C_{4}}{n_{k}}}\big{)}n_{k}}=o(1),$
where $R_{3}$ is as in Condition 1. In addition,
$\frac{\sqrt{n_{0}+n_{\mathcal{A}}}}{\min_{1\leq k\leq K}n_{k}}=o(1),\quad\frac{\sqrt{n_{0}+n_{\mathcal{A}}}}{n_{0}}=o(1).$
Remark 1.
These conditions are typically satisfied in practice as they are not overly restrictive. Condition 1 requires that covariates and covariance matrices are bounded in a specific way, which is standard in the transfer learning literature and generally holds in real-world scenarios (Cai et al. 2024). In particular, this assumption is common when dealing with high-dimensional data where regularization is necessary (Li et al. 2022). Condition 2 assumes that the number of samples is significantly larger than the number of sources $K$, which is reasonable since $K$ is usually fixed in practical settings, and we often have sufficient source data compared to target data. In practice, Condition 2 may be slightly violated if there exists a source $k$ with a relatively small $n_{k}$, such that $\sqrt{n_{0}+n_{\mathcal{A}}}/n_{k}$ is not $o(1)$. In such cases, the $k$th source can simply be excluded from Step 1 of the WaTL algorithm.
Theorem 1.
Suppose Conditions 1 and 2 hold. Then, for the WaTL algorithm, it holds for a fixed $x\in\mathbb{R}^{p}$ that
$\|\widehat{f}(x)-f(x)\|_{2}=O_{p}\Big{(}\frac{\sum_{k=0}^{K}\sqrt{n_{k}}}{n_{0}+n_{\mathcal{A}}}+(n_{0}+n_{\mathcal{A}})^{-1/2}\Big{)},$
where $f(x)=(n_{0}+n_{\mathcal{A}})^{-1}\sum_{k=0}^{K}n_{k}f^{(k)}(x)$.
The proof of Theorem 1 involves a detailed analysis of the sample covariance matrix and leverages M-estimation techniques within the framework of empirical process theory. The result extends beyond responses in the Wasserstein space, applying to other metric spaces that meet mild regularity conditions. Consequently, Theorem 1 provides a versatile framework that can be applied to transfer learning in regression models with responses such as networks (Zhou & Müller 2022), symmetric positive-definite matrices (Pennec et al. 2006), or trees (Billera et al. 2001). The following theorem establishes the convergence rate for the estimated regression function $\widehat{m}_{G}^{(0)}(x)$ in the WaTL algorithm.
Theorem 2.
Assume Conditions 1 and 2 hold and the regularization parameter satisfies $\lambda\asymp n_{0}^{-1/2+\epsilon}$ for some $\epsilon>0$. Then, for the WaTL algorithm and a fixed $x\in\mathbb{R}^{p}$, it holds that
$d_{\mathcal{W}}^{2}(\widehat{m}_{G}^{(0)}(x),m_{G}^{(0)}(x))=O_{p}\Big{(}n_{0}^{-1/2+\epsilon}\big{(}\psi+\frac{\sum_{k=0}^{K}\sqrt{n_{k}}}{n_{0}+n_{\mathcal{A}}}+(n_{0}+n_{\mathcal{A}})^{-1/2}\big{)}\Big{)},$
where $\psi=\max_{1\leq k\leq K}\|f^{(0)}(x)-f^{(k)}(x)\|_{2}$ quantifies the maximum discrepancy between the target and source.
Theorem 2 can be contrasted with the convergence rate of global Fréchet regression (Petersen & Müller 2019) applied solely to the target data, where the rate for regression with distributional outputs is $d_{\mathcal{W}}(\widehat{m}_{G}^{(0)}(x),m_{G}^{(0)}(x))=O_{p}(n_{0}^{-1/2})$. The WaTL algorithm achieves a faster convergence rate when there are sufficient source data and the auxiliary sources are informative enough, satisfying $\psi\lesssim n_{0}^{-1/2-\epsilon}$. This result highlights that knowledge transfer from auxiliary samples can significantly enhance the learning performance of the target model, provided the auxiliary sources are closely aligned with the target data.
For the AWaTL algorithm, we require the following condition.
Condition 3.
The regularization parameter satisfies $\lambda\asymp n_{0}^{-1/2+\epsilon}$ for some $\epsilon>0$ and for some $\epsilon^{\prime}>\epsilon$, there exists a non-empty subset $\mathcal{A}\subset\{1,\ldots,K\}$ such that
$\frac{n_{*}^{-1/2}+n_{0}^{-1/2}}{\min_{k\in\mathcal{A}^{C}}\psi_{k}}=o(1),\quad\frac{\max_{k\in\mathcal{A}}\psi_{k}}{n_{*}^{-1/2}+n_{0}^{-1/2-\epsilon^{\prime}}}=O(1),$
[…]
Remark 2.
We allow $\mathcal{A}$ to be the whole set $\{1,\ldots,K\}$, in which case the condition becomes
$\frac{\max_{1\leq k\leq K}\psi_{k}}{n_{*}^{-1/2}+n_{0}^{-1/2-\epsilon^{\prime}}}=O(1).$
Besides, if $\mathcal{A}$ exists, then it is unique since
$\{1\leq k\leq K:\frac{n_{*}^{-1/2}+n_{0}^{-1/2}}{\psi_{k}}=o(1)\}\cap\{1\leq k\leq K:\frac{\psi_{k}}{n_{*}^{-1/2}+n_{0}^{-1/2-\epsilon^{\prime}}}=O(1)\}=\emptyset.$
Condition 3 ensures that the source datasets can be effectively partitioned into two groups: one that is informative and another that is sufficiently different from the target data. This separation guarantees that the informative set can be accurately identified. Under this condition, setting the number of informative sources as $L=|\mathcal{A}|$, we establish the following rate of convergence for the AWaTL algorithm. In practice, $L$ can be decided by cross-validation.
Theorem 3.
Under Condition 3 and the conditions of Theorem 2, for the AWaTL algorithm with a fixed number of sources $K$ we have $P(\widehat{\mathcal{A}}=\mathcal{A})\rightarrow 1$ and
$d_{\mathcal{W}}^{2}(\widehat{m}_{G}^{(0)}(x),m_{G}^{(0)}(x))=O_{p}\Big{(}n_{0}^{-1/2+\epsilon}\big{(}\max_{k\in\mathcal{A}}\psi_{k}+\frac{\sum_{k\in\mathcal{A}\cup\{0\}}\sqrt{n_{k}}}{\sum_{k\in\mathcal{A}\cup\{0\}}n_{k}}+(\sum_{k\in\mathcal{A}\cup\{0\}}n_{k})^{-1/2}\big{)}\Big{)}.$
The convergence rate of the AWaTL algorithm simplifies to $n_{0}^{-1-\epsilon^{\prime}+\epsilon}$ when the informative source data is sufficiently large. This rate surpasses that of global Fréchet regression (Petersen & Müller 2019) applied solely to the target data, offering a theoretical guarantee that AWaTL effectively mitigates negative transfer by selectively integrating relevant auxiliary information.
5 Numerical Experiments
In this section, we evaluate the performance of the proposed WaTL algorithm, alongside two baseline approaches: the global Fréchet regression using only target data (Only Target) and using only source data (Only Source). Consider $K=5$ source sites. The data is generated as follows. For the target population, we sample $X^{(0)}\sim{\rm U}(0,1)$ and generate the response distribution, represented by its quantile function, as
$F^{-1}_{\nu^{(0)}}(u)=w^{(0)}(1-u)u+(1-X^{(0)})u+X^{(0)}F^{-1}_{Z^{(0)}}(u),\quad u\in(0,1),$
where $Z^{(0)}\sim N(0.5,1)|_{(0,1)}$ and $w^{(0)}\sim N(0,1)|_{(-0.5,0.5)}$. Here, $N(\mu,\sigma^{2})|_{(a,b)}$ denotes a normal distribution with mean $\mu$ and variance $\sigma^{2}$, truncated to the interval $(a,b)$. For source populations, we define $\psi_{k}=0.1k$ for $k=1,\ldots,K$ and generate $X^{(k)}\sim{\rm U}(0,1)$. The corresponding response distribution is generated as
$F^{-1}_{\nu^{(k)}}(u)=w^{(k)}(1-u)u+(1-X^{(k)})u+X^{(k)}F^{-1}_{Z^{(k)}}(u),\quad u\in(0,1),$
where $Z^{(k)}\sim N(0.5,1-\psi_{k})|_{(0,1)}$ and $w^{(k)}\sim N(0,1)|_{(-0.5,0.5)}$. Consequently, for each predictor $x$, the true regression function is $m_{G}^{(k)}(x)=(1-x)u+xF^{-1}_{Z^{(k)}}(u)$ for $k=0,1,\ldots,K$.
We vary the target sample size $n_{0}$ from $200$ to $800$, while the source sample size is set as $n_{k}=k\tau$, where $\tau\in\{100,200\}$ and $k=1,\ldots,K$. The regularization parameter $\lambda$ in Algorithm 1 is selected via five-fold cross-validation, varying from 00 to $3$ in increments of $0.1$. To evaluate performance, we sample $100$ predictors uniformly from the target distribution. Using Algorithm 1, we compute $\widehat{m}_{G}^{(0)}(x)$ and compare it with the corresponding estimates obtained from global Fréchet regression using only target or source data. Performance is assessed using the root mean squared prediction risk $\mathrm{RMSPR}=\sqrt{\frac{1}{100}\sum_{i=1}^{100}d_{\mathcal{W}}^{2}\big{(}\widehat{m}_{G}^{(0)}(x_{i}),{m}_{G}^{(0)}(x_{i})\big{)}}$, where $x_{i}$ denotes the sampled predictor, $\widehat{m}_{G}^{(0)}(x_{i})$ is the estimated function, and ${m}_{G}^{(0)}(x_{i})$ represents the ground truth. To ensure robustness, we repeat the simulation $50$ times and report the average RMSPR, as shown in Figure 1 (a).
As shown in Figure 1 (a), WaTL consistently outperforms global Fréchet regression trained solely on target or source data. When the target sample size $n_{0}$ is small, the Only Target method exhibits a high RMSPR due to the instability of models trained on limited data. In contrast, WaTL significantly reduces RMSPR by effectively incorporating auxiliary information from the source domain. As $n_{0}$ increases, the performance of Only Target improves and gradually approaches that of WaTL, which is expected as larger sample sizes lead to more stable and accurate estimators. Nevertheless, WaTL maintains a consistent advantage across all $n_{0}$, suggesting that it leverages complementary information from the source. The performance of the Only Source estimator remains nearly unchanged across different $n_{0}$ values, as it does not benefit from additional target data.
Comparing the two panels of Figure 1 (a), we also observe that WaTL improves as the source sample size increases, confirming its ability to effectively integrate source information. This demonstrates the benefit of multi-source transfer learning, where WaTL balances knowledge from both target and source domains to achieve improved prediction.
To better understand when negative transfer may occur, we conduct an ablation study with $K=1$ source and vary the similarity parameter $\psi_{1}$ from $0.01$ to $1$ in increments of $0.01$, with $n_{0}=100$ and $n_{1}=200$. Our method outperforms the Only Target approach when $\psi_{1}<0.9$, while for $\psi_{1}\geq 0.9$, the Only Target method becomes preferable. This confirms that negative transfer arises when the source is too dissimilar to the target.
We further evaluate the effectiveness of AWaTL in selecting informative sources. In this experiment, we set $L=2$ and $n_{k}=100$ for $k=0,\ldots,5$. The similarity parameters are set as $\psi_{k}=0.1$ for $k=1,2$ (informative sources), and $\psi_{k}$ from $0.2$ to $1$ in increments of $0.1$ for $k=3,4,5$ (uninformative sources). We repeat each setting 100 times and report the selection rates in Figure 1 (b). The results show that AWaTL successfully identifies informative sources, with selection rates for sources $1$ and $2$ rapidly increasing and reaching perfect accuracy when $\psi>0.6$. This highlights the robustness of AWaTL in distinguishing useful sources under varying similarity levels.
Figure 1: (a) Root mean squared prediction risk (RMSPR) of WaTL, only Source, and Only Target methods under varying target sample sizes, with source sample sizes $\tau=100$ (left) and $\tau=200$ (right); (b) Selection rate of each source site as $\psi$ increases.
6 Real-world Applications
We evaluate the WaTL algorithm using data from the National Health and Nutrition Examination Survey (NHANES) 2005–2006111https://wwwn.cdc.gov/nchs/nhanes/ContinuousNhanes/Default.aspx?BeginYear=2005, focusing on modeling the distribution of physical activity intensity. NHANES is a large-scale health survey in the United States that combines interviews with physical examinations to assess the health and nutrition of both adults and children. The dataset includes extensive demographic, socioeconomic, dietary, and medical assessments, providing a comprehensive resource for health-related research.
During the 2005–2006 NHANES cycle, participants aged 6 and older wore an ActiGraph 7164 accelerometer on their right hip for seven days, recording physical activity intensity in 1-minute epochs. Participants were instructed to remove the device during water-based activities and sleep. The device measured counts per minute (CPM), ranging from 00 to $32767$, capturing variations in activity levels throughout the monitoring period.
Since female and male participants exhibit distinct physical activity patterns (Feraco et al. 2024), we analyze them separately. Physical activity intensity is influenced by multiple factors, and we consider body mass index (BMI) and age as key predictors (Klein & Becker 2012, Cleven et al. 2023). To accommodate potential nonlinear relationships, we implement local Fréchet regression within the WaTL algorithm, treating the distribution of physical activity intensity as the response and BMI and age as predictors.
Following the data preprocessing steps in Lin et al. (2023), we remove unreliable observations per NHANES protocols. For each participant, we exclude activity counts above 1000 CPM or equal to zero (as zeros may correspond to various low-activity states such as sleep or swimming). Participants with fewer than $100$ valid observations or missing BMI, age, or gender information are also excluded. The remaining activity counts over seven days are concatenated to form the distribution of each participant’s activity intensity.
To evaluate the WaTL, we set White, Mexican Americans, and other Hispanic individuals as sources and Black as target. For females, the source data include $1308$ White people, $884$ Mexican Americans, and $108$ Other Hispanic individuals. For males, the source data include $1232$ White people, $805$ Mexican Americans, and $92$ Other Hispanic individuals. We set $200$ Black participants as the target data for both genders. During evaluating, we perform five-fold cross-validation, using four folds for training and one for testing, cycling through each fold. For comparison, we also apply local Fréchet regression using only the target data. The results, summarized in Figure 2 (a), show that WaTL improves performance over local Fréchet regression for both females and males, demonstrating its ability to leverage information from other participants to enhance modeling for black participants.
Figure 2: (a) Root mean squared prediction risk (RMSPR) of WaTL and Only Target methods for females and males, evaluated using five-fold cross-validation; (b) Cumulative distribution function of physical activity levels for one selected female (left) and one selected male (right), along with estimates from WaTL and Only Target methods.
[…]
7 Conclusion
We introduce Wasserstein transfer learning and its adaptive variant for scenarios where the informative set is unknown, addressing the challenges posed by the lack of linear operations in the Wasserstein space. By leveraging the Wasserstein metric, the proposed algorithm accounts for the non-Euclidean structure of distributional outputs, ensuring compatibility with the geometry of the Wasserstein space. Supported by rigorous theoretical guarantees, the framework demonstrates improved estimation performance compared to methods that rely solely on target data.
This paper focuses on univariate distributions, which arise frequently in real-world applications such as the NHANES and human mortality studies discussed in Section 6. The proposed framework, however, is not limited to the univariate case and can be extended to multivariate distributions by incorporating metrics such as the Sinkhorn (Cuturi 2013) or sliced Wasserstein distance (Bonneel et al. 2015). Extending the methodology to higher dimensions presents an exciting future direction and would involve addressing both computational and theoretical challenges specific to high-dimensional Wasserstein spaces.
Our theoretical analysis assumes that each probability distribution is fully observed, consistent with much of the existing literature on distributional data analysis (Petersen & Müller 2019, Chen et al. 2023). However, this assumption could be unrealistic in practical applications, where only independent samples from the underlying distributions are available. This limitation can be addressed by replacing the unobservable distributions with empirical measures constructed from these samples (Zhou & Müller 2024). The derived convergence rates can be extended to incorporate an additional term reflecting the number of independent samples per distribution, providing a straightforward way to account for this practical consideration.
[…]
We further assess WaTL using the age-at-death distributions from 162 countries in 2015, compiled from the United Nations Databases [^2] https://data.un.org/ and the UN World Population Prospects 2022333https://population.un.org/wpp/Download. The dataset provides country-level age-specific death counts, which we convert into smooth age-at-death densities using local linear smoothing; see Figure 3 (a) for an illustration. For this analysis, we define the 24 developed countries as the target site and the remaining 138 developing countries as the source site.
Given the limited sample size, we input the entire dataset into the model without performing cross-validation. Country-level age-at-death distributions are shaped by a range of factors, including economic conditions, healthcare systems, and social and environmental contexts. Among these, we select population density and gross domestic product (GDP) per capita as key predictive covariates. The results, shown in Figure 3 (b), demonstrate that the proposed WaTL method successfully transfers knowledge across these demographically diverse regions, effectively addressing the socioeconomic disparities between the source and target domains.
Figure 3: (a) Age-at-death densities of developed and developing countries; (b) Root mean squared prediction risk (RMSPR) of WaTL and Only Target methods for human mortality data.

### v2

@@
 However, traditional transfer learning approaches often focus on scalar or multivariate data within Euclidean spaces, limiting their applicability to complex data structures such as probability distributions.
-To address this, we introduce a novel framework for transfer learning in regression models, where outputs are probability distributions residing in the Wasserstein space.
+To address this limitation, we introduce a novel transfer learning framework for regression models whose outputs are probability distributions residing in the Wasserstein space.
 When the informative subset of transferable source domains is known, we propose an estimator with provable asymptotic convergence rates, quantifying the impact of domain similarity on transfer efficiency.
@@
 The proposed methods are supported by rigorous theoretical analysis and are validated through extensive simulations and real-world applications.
+The code is available at https://github.com/h7nian/WaTL.
 Keywords: Data Aggregation, Domain Adaptation, Multitask Learning, non-Euclidean Output, Optimal Transport
@@
 2012, Choi et al.
-2017, Zhou et al.
-2024).
+2017, Zhou, Liu, Li & Cai 2024, Zhou, Li, Cai & Liu 2024).
 Recent theoretical developments have focused on regression in Euclidean settings, including high-dimensional linear (Li et al.
@@
 2.1 Notations
-[…]
+Let $L^{2}(0,1)$ be the space of square-integrable functions over the interval $(0,1)$, with the associated $L^{2}$ norm and metric denoted by $\|\cdot\|_{2}$ and $d_{L^{2}}$, respectively.
+To be specific, $\|g\|_{2}=(\int_{0}^{1}g^{2}(z)dz)^{1/2}\text{ and }d_{L^{2}}(g_{1},g_{2})=\|g_{1}-g_{2}\|_{2}$.
+For a vector $Z$, $\|Z\|$ denotes the Euclidean norm.
+Given a matrix $\Sigma$, we define its spectrum as the set of its singular values.
+For a sub-Gaussian random vector $X$, we define the sub-Gaussian norm as $\|X\|_{\Psi_{2}}:=\sup_{\|v\|=1}\inf\left\{t>0:E(e^{\langle X,v\rangle^{2}/t^{2}})\leq 2\right\}.$
 We write $a_{n}\lesssim b_{n}$ if there exists a positive constant $C$ such that $a_{n}\leq Cb_{n}$ when $n$ is large enough and $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$.
@@
 2.2 Wasserstein Space
-[…]
-[…]
+Let $\mathcal{W}$ denote the space of probability distributions on $\mathbb{R}$ with finite second moments, equipped with the 2-Wasserstein, or simply Wasserstein, metric.
+For two distributions $\mu_{1},\mu_{2}\in\mathcal{W}$, the Wasserstein metric is given by $d_{\mathcal{W}}^{2}(\mu_{1},\mu_{2})=\inf_{\pi\in\Pi(\mu_{1},\mu_{2})}\int_{\mathbb{R}\times\mathbb{R}}|s-t|^{2}\,d\pi(s,t)$, where $\Pi(\mu_{1},\mu_{2})$ denotes the set of all joint distributions with marginals $\mu_{1}$ and $\mu_{2}$ (Kantorovich 1942).
+For a probability measure $\mu\in\mathcal{W}$ with cumulative distribution function $F_{\mu}$, we define the quantile function $F_{\mu}^{-1}$ as the left-continuous inverse of $F_{\mu}$, such that $F_{\mu}^{-1}(u)=\inf\{t\in\mathbb{R}|F_{\mu}(t)\geq u\}$, for $u\in(0,1)$.
+It has been established (Villani 2009) that the Wasserstein metric can be expressed as the $L^{2}$ metric between quantile functions:
+$d_{\mathcal{W}}^{2}(\mu_{1},\mu_{2})=\int_{0}^{1}\left\{F_{\mu_{1}}^{-1}(u)-F_{\mu_{2}}^{-1}(u)\right\}^{2}du.$ (1)
 The space $\mathcal{W}$, endowed with the Wasserstein metric, forms a complete and separable metric space, commonly known as the Wasserstein space (Villani 2009).
@@
 We consider a transfer learning problem where target data $\{(X_{i}^{(0)},\nu_{i}^{(0)})\}_{i=1}^{n_{0}}$ are sampled independently from the target population $(X^{(0)},\nu^{(0)})\sim\mathcal{F}_{0}$, and source data $\{(X_{i}^{(k)},\nu_{i}^{(k)})\}_{i=1}^{n_{k}}$ are sampled independently from the source population $(X^{(k)},\nu^{(k)})\sim\mathcal{F}_{k}$, for $k=1,\ldots,K$.
-The goal is to estimate the target model using both the target data and the source data from $K$ related studies.
+The goal is to estimate the target model using both the target data and source data from $K$ related studies.
 For $k=0,\ldots,K$, assume $X^{(k)}$ has mean $\theta_{k}$ and covariance $\Sigma_{k}$, with $\Sigma_{k}$ positive definite.
-Define the empirical mean and covariance of $\{X_{i}^{(k)}\}_{i=1}^{n_{k}}$ as $\overline{X}_{k}=n_{k}^{-1}\sum_{i=1}^{n_{k}}X_{i}^{(k)}$ and $\widehat{\Sigma}_{k}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}(X_{i}^{(k)}-\overline{X}_{k})(X_{i}^{(k)}-\overline{X}_{k})^{\mathrm{\scriptscriptstyle T}}.$ For a fixed $x\in\mathbb{R}^{p}$, the weight function is $s_{G}^{(k)}(x)=1+(X^{(k)}-\theta_{k})^{\mathrm{\scriptscriptstyle T}}\Sigma_{k}^{-1}(x-\theta_{k}),$ with the sample version $s_{iG}^{(k)}(x)=1+(X_{i}^{(k)}-\overline{X}_{k})^{\mathrm{\scriptscriptstyle T}}\widehat{\Sigma}_{k}^{-1}(x-\overline{X}_{k}).$ The target regression function for a given $x\in\mathbb{R}^{p}$ is then $m_{G}^{(0)}(x)=\operatorname*{arg\,min}_{\mu\in\mathcal{W}}E\{s_{G}^{(0)}(x)d_{\mathcal{W}}^{2}(\mu,\nu^{(0)})\}.$ In the following, we present details on transfer learning for global Fréchet regression, where the key difference in the local Fréchet regression setting is the use of a different weight function.
-The technical details for transfer learning in local Fréchet regression are therefore deferred to Appendix D.
+Define the empirical mean and covariance of $\{X_{i}^{(k)}\}_{i=1}^{n_{k}}$ as $\overline{X}_{k}=n_{k}^{-1}\sum_{i=1}^{n_{k}}X_{i}^{(k)}$ and $\widehat{\Sigma}_{k}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}(X_{i}^{(k)}-\overline{X}_{k})(X_{i}^{(k)}-\overline{X}_{k})^{\mathrm{\scriptscriptstyle T}}.$ For a fixed $x\in\mathbb{R}^{p}$, the weight function is $s_{G}^{(k)}(x)=1+(X^{(k)}-\theta_{k})^{\mathrm{\scriptscriptstyle T}}\Sigma_{k}^{-1}(x-\theta_{k}),$ with the sample version $s_{iG}^{(k)}(x)=1+(X_{i}^{(k)}-\overline{X}_{k})^{\mathrm{\scriptscriptstyle T}}\widehat{\Sigma}_{k}^{-1}(x-\overline{X}_{k}).$ The target regression function for a given $x\in\mathbb{R}^{p}$ is then $m_{G}^{(0)}(x)=\operatorname*{arg\,min}_{\mu\in\mathcal{W}}E\{s_{G}^{(0)}(x)d_{\mathcal{W}}^{2}(\nu^{(0)},\mu)\}.$ In the following, we present details on transfer learning for global Fréchet regression, where the key difference in the local Fréchet regression setting is the use of a different weight function.
+The technical details for transfer learning in local Fréchet regression are therefore deferred to Appendix 11.
 The set of informative auxiliary samples (informative set) consists of sources sufficiently similar to the target data.
@@
 3.2 Wasserstein Transfer Learning
-We propose the Wasserstein Transfer Learning (WaTL) algorithm, which combines information from source datasets under the assumption that all source data are “informative enough.” This assumption implies that the discrepancies between the source and target are small enough to enhance estimation compared to using only the target.
+We propose the Wasserstein Transfer Learning (WaTL) algorithm, which combines information from source datasets under the assumption that all source data are informative enough.
+This assumption implies that the discrepancies between the source and target are small enough to enhance estimation compared to using only the target.
 When this condition is met for all source datasets, the informative set is given by $\mathcal{A}_{\psi}=\{1,\ldots,K\}$.
@@
 3: Projection to Wasserstein space: $\widehat{m}^{(0)}_{G}(x)=\operatorname*{arg\,min}_{\mu\in\mathcal{W}}\big\|F^{-1}_{\mu}-\widehat{f}_{0}(x)\big\|_{2}.$
-In Step 1, the initial estimate $\widehat{f}$ aggregates information from both the target and source, weighted by their respective sample numbers.
+In Step 1, the initial estimate $\widehat{f}$ aggregates information from both the target and source, weighted by their respective sample sizes.
 While this step incorporates valuable auxiliary information, the resulting estimate may be biased due to distributional differences between the target and source populations.
@@
 Theoretical guidelines for selecting $\lambda$ are provided in Theorem 2.
-The final step projects the corrected estimate $\widehat{f}_{0}$ onto the Wasserstein space, ensuring the output $\widehat{m}^{(0)}_{G}(x)$ satisfies the intrinsic geometry of $\mathcal{W}$.
-Such a projection exists and is unique as $\mathcal{W}$ is a closed and convex subset of $L^{2}(0,1)$.
+The final step projects the corrected estimate $\widehat{f}_{0}$ onto the Wasserstein space, ensuring the output $\widehat{m}^{(0)}_{G}(x)$ respects the intrinsic geometry of $\mathcal{W}$.
+This projection exists and is unique because $\mathcal{W}$ is a closed and convex subset of $L^{2}(0,1)$.
 3.3 Adaptive Selection of Informative Sources
@@
 Algorithm 2 Adaptive Wasserstein Transfer Learning (AWaTL)
-0:  Target and source data $\{(x_{i}^{(0)},\nu_{i}^{(0)})\}_{i=1}^{n_{0}}\cup\big{(}\cup_{1\leq k\leq K}\{(x_{i}^{(k)},\nu_{i}^{(k)})\}_{i=1}^{n_{k}}\big{)}$, regularization parameter $\lambda$, number of informative sources $L$, and query point $x\in\mathbb{R}^{p}$.
+0: Target and source data $\{(x_{i}^{(0)},\nu_{i}^{(0)})\}_{i=1}^{n_{0}}\cup\big(\cup_{1\leq k\leq K}\{(x_{i}^{(k)},\nu_{i}^{(k)})\}_{i=1}^{n_{k}}\big)$, regularization parameter $\lambda$, prespecified number of informative sources $L$, and query point $x\in\mathbb{R}^{p}$.
 0: Target estimator $\widehat{m}_{G}^{(0)}(x)$.
-1:  Compute discrepancy scores.
-For each source dataset $k=1,\ldots,K$, compute the empirical discrepancy: $\widehat{\psi}_{k}=\|\widehat{f}^{(0)}(x)-\widehat{f}^{(k)}(x)\|_{2},$ where $\widehat{f}^{(k)}(x)=n_{k}^{-1}\sum_{i=1}^{n_{k}}s_{iG}^{(k)}(x)F^{-1}_{\nu_{i}^{(k)}}$.
-Construct the adaptive informative set by selecting the $L$ smallest discrepancy scores $\widehat{\mathcal{A}}=\{1\leq k\leq K:\widehat{\psi}_{k}\text{ is among the smallest }L\text{ values}\}.$
-2:  Weighted auxiliary estimator: $\widehat{f}(x)=\frac{1}{\sum_{k\in\widehat{\mathcal{A}}\cup\{0\}}n_{k}}\sum_{k\in\widehat{\mathcal{A}}\cup\{0\}}n_{k}\widehat{f}^{(k)}(x).$
+[…]
+[…]
 3: Bias correction using target data: $\widehat{f}_{0}(x)=\operatorname*{arg\,min}_{g\in L^{2}(0,1)}\frac{1}{n_{0}}\sum_{i=1}^{n_{0}}s_{iG}^{(0)}(x)\|F^{-1}_{\nu_{i}^{(0)}}-g\|_{2}^{2}+\lambda\|g-\widehat{f}(x)\|_{2}.$
@@
 Condition 1.
-For $k=0,\ldots,K$, the covariate $X^{(k)}$ is subguassian with $\|X^{(k)}\|_{\Psi_{2}}\in[\sigma_{1},\sigma_{2}]$, the mean vector satisfies $\|\theta_{k}\|\leq R_{1}$, and the spectrum of the covariance matrix $\Sigma_{k}$ lies within the interval $[R_{2},R_{3}]$.
+For $k=0,\ldots,K$, the covariate $X^{(k)}$ is sub-Gaussian with $\|X^{(k)}\|_{\Psi_{2}}\in[\sigma_{1},\sigma_{2}]$, the mean vector satisfies $\|\theta_{k}\|\leq R_{1}$, and the spectrum of the covariance matrix $\Sigma_{k}$ lies within the interval $R_{2},R_{3}]$.
 Moreover, $\nu^{(k)}$ is supported on a bounded interval.
 Lemma 1.
-Let $\widehat{f}^{(k)}(x)=n_{k}^{-1}\sum_{i=1}^{n_{k}}s_{iG}^{(k)}(x)F_{\nu_{i}^{(k)}}^{-1}$ and its population counterpart $f^{(k)}(x)=E\{s_{G}^{(k)}(x)F_{\nu^{(k)}}^{-1}\}$ for $k=0,\ldots,K$.
+Let $\widehat{f}^{(k)}(x)=n_{k}^{-1}\sum_{i=1}^{n_{k}}s_{iG}^{(k)}(x)F_{\nu_{i}^{(k)}}^{-1}$ and its population counterpart be defined as $f^{(k)}(x)=E\{s_{G}^{(k)}(x)F_{\nu^{(k)}}^{-1}\}$ for $k=0,\ldots,K$.
 Then under Condition 1, $\|\widehat{f}^{(k)}(x)-f^{(k)}(x)\|_{2}=O_{p}(n_{k}^{-1/2})$.
@@
 where $\psi=\max_{1\leq k\leq K}\|f^{(0)}(x)-f^{(k)}(x)\|_{2}$ quantifies the maximum discrepancy between the target and source.
-Theorem 2 can be contrasted with the convergence rate of global Fréchet regression (Petersen & Müller 2019) applied solely to the target data, where the rate for regression with distributional outputs is $d_{\mathcal{W}}(\widehat{m}_{G}^{(0)}(x),m_{G}^{(0)}(x))=O_{p}(n_{0}^{-1/2})$.
+Theorem 2 can be compared to the convergence rate of global Fréchet regression (Petersen & Müller 2019) applied solely to the target data, for which the rate is $d_{\mathcal{W}}(\widehat{m}_{G}^{(0)}(x),m_{G}^{(0)}(x))=O_{p}(n_{0}^{-1/2})$.
 The WaTL algorithm achieves a faster convergence rate when there are sufficient source data and the auxiliary sources are informative enough, satisfying $\psi\lesssim n_{0}^{-1/2-\epsilon}$.
@@
 $\{1\leq k\leq K:\frac{n_{*}^{-1/2}+n_{0}^{-1/2}}{\psi_{k}}=o(1)\}\cap\{1\leq k\leq K:\frac{\psi_{k}}{n_{*}^{-1/2}+n_{0}^{-1/2-\epsilon^{\prime}}}=O(1)\}=\emptyset.$
-Condition 3 ensures that the source datasets can be effectively partitioned into two groups: one that is informative and another that is sufficiently different from the target data.
+Condition 3 ensures that the source datasets can be effectively partitioned into two groups: informative ones and those sufficiently different from the target data.
 This separation guarantees that the informative set can be accurately identified.
-Under this condition, setting the number of informative sources as $L=|\mathcal{A}|$, we establish the following rate of convergence for the AWaTL algorithm.
+Under this condition, by setting the number of informative sources to $L=|\mathcal{A}|$, we establish the following rate of convergence for the AWaTL algorithm.
 In practice, $L$ can be decided by cross-validation.
@@
 This rate surpasses that of global Fréchet regression (Petersen & Müller 2019) applied solely to the target data, offering a theoretical guarantee that AWaTL effectively mitigates negative transfer by selectively integrating relevant auxiliary information.
+While the above analysis assumes that each probability distribution is fully observed, an assumption commonly adopted in the distributional data literature (Petersen & Müller 2019, Chen et al.
+2023), real-world applications often provide only independent samples drawn from underlying distributions.
+In such cases, this limitation can be overcome by replacing unobservable distributions with their empirical counterparts, constructed from sample observations (Zhou & Müller 2024).
+Additional details and theoretical justification for this extension are provided in Appendix 12.
 5 Numerical Experiments
@@
 Consider $K=5$ source sites.
-The data is generated as follows.
+The data are generated as follows.
 For the target population, we sample $X^{(0)}\sim{\rm U}(0,1)$ and generate the response distribution, represented by its quantile function, as
@@
 Here, $N(\mu,\sigma^{2})|_{(a,b)}$ denotes a normal distribution with mean $\mu$ and variance $\sigma^{2}$, truncated to the interval $(a,b)$.
-For source populations, we define $\psi_{k}=0.1k$ for $k=1,\ldots,K$ and generate $X^{(k)}\sim{\rm U}(0,1)$.
+For source populations, we define $\psi_{k}=0.1k$ for $k=1,\ldots,K$, and generate $X^{(k)}\sim{\rm U}(0,1)$.
 The corresponding response distribution is generated as
@@
 where $Z^{(k)}\sim N(0.5,1-\psi_{k})|_{(0,1)}$ and $w^{(k)}\sim N(0,1)|_{(-0.5,0.5)}$.
-Consequently, for each predictor $x$, the true regression function is $m_{G}^{(k)}(x)=(1-x)u+xF^{-1}_{Z^{(k)}}(u)$ for $k=0,1,\ldots,K$.
+Consequently, for each predictor $x$, the true regression function is $m_{G}^{(k)}(x)=(1-x)u+xF^{-1}_{Z^{(k)}}(u)$, for $k=0,1,\ldots,K$.
 We vary the target sample size $n_{0}$ from $200$ to $800$, while the source sample size is set as $n_{k}=k\tau$, where $\tau\in\{100,200\}$ and $k=1,\ldots,K$.
-The regularization parameter $\lambda$ in Algorithm 1 is selected via five-fold cross-validation, varying from 00 to $3$ in increments of $0.1$.
+The regularization parameter $\lambda$ in Algorithm 1 is selected via five-fold cross-validation, ranging from 0 to $3$ in increments of $0.1$.
 To evaluate performance, we sample $100$ predictors uniformly from the target distribution.
@@
 The performance of the Only Source estimator remains nearly unchanged across different $n_{0}$ values, as it does not benefit from additional target data.
-Comparing the two panels of Figure 1 (a), we also observe that WaTL improves as the source sample size increases, confirming its ability to effectively integrate source information.
+Comparing the two panels of Figure 1(a), we also observe that WaTL improves as the source sample size increases, confirming its ability to effectively integrate information from source domains.
 This demonstrates the benefit of multi-source transfer learning, where WaTL balances knowledge from both target and source domains to achieve improved prediction.
@@
 In this experiment, we set $L=2$ and $n_{k}=100$ for $k=0,\ldots,5$.
-The similarity parameters are set as $\psi_{k}=0.1$ for $k=1,2$ (informative sources), and $\psi_{k}$ from $0.2$ to $1$ in increments of $0.1$ for $k=3,4,5$ (uninformative sources).
-We repeat each setting 100 times and report the selection rates in Figure 1 (b).
-The results show that AWaTL successfully identifies informative sources, with selection rates for sources $1$ and $2$ rapidly increasing and reaching perfect accuracy when $\psi>0.6$.
-This highlights the robustness of AWaTL in distinguishing useful sources under varying similarity levels.
+The similarity parameters are specified as $\psi_{k}=0.1$ for $k=1,2$ (informative sources), and $\psi_{k}=\psi$, increasing from $0.2$ to $1$ in increments of $0.1$, for $k=3,4,5$ (uninformative sources).
+Each configuration is repeated $100$ times, and the corresponding selection rates are reported in Figure 1(b).
+The results show that AWaTL successfully identifies informative sources, with selection rates for sources $1$ and $2$ rapidly increasing and reaching perfect accuracy once $\psi>0.6$.
+These findings demonstrate the robustness of AWaTL in distinguishing useful sources under varying similarity levels.
 Figure 1: (a) Root mean squared prediction risk (RMSPR) of WaTL, only Source, and Only Target methods under varying target sample sizes, with source sample sizes $\tau=100$ (left) and $\tau=200$ (right); (b) Selection rate of each source site as $\psi$ increases.
@@
 Participants were instructed to remove the device during water-based activities and sleep.
-The device measured counts per minute (CPM), ranging from 00 to $32767$, capturing variations in activity levels throughout the monitoring period.
+The device measured counts per minute (CPM), ranging from 0 to $32767$, capturing variations in activity levels throughout the monitoring period.
 Since female and male participants exhibit distinct physical activity patterns (Feraco et al.
@@
 The remaining activity counts over seven days are concatenated to form the distribution of each participant’s activity intensity.
-To evaluate the WaTL, we set White, Mexican Americans, and other Hispanic individuals as sources and Black as target.
+To evaluate the WaTL, we set White, Mexican Americans, and other Hispanic individuals as sources and Black as the target.
 For females, the source data include $1308$ White people, $884$ Mexican Americans, and $108$ Other Hispanic individuals.
-For males, the source data include $1232$ White people, $805$ Mexican Americans, and $92$ Other Hispanic individuals.
+For males, the source data include $1232$ White participants, $805$ Mexican Americans, and $92$ Other Hispanic individuals.
 We set $200$ Black participants as the target data for both genders.
@@
 For comparison, we also apply local Fréchet regression using only the target data.
-The results, summarized in Figure 2 (a), show that WaTL improves performance over local Fréchet regression for both females and males, demonstrating its ability to leverage information from other participants to enhance modeling for black participants.
+The results, summarized in Figure 2(a), show that WaTL improves performance over local Fréchet regression for both females and males, demonstrating its ability to leverage information from other demographic groups to enhance modeling for Black participants.
 Figure 2: (a) Root mean squared prediction risk (RMSPR) of WaTL and Only Target methods for females and males, evaluated using five-fold cross-validation; (b) Cumulative distribution function of physical activity levels for one selected female (left) and one selected male (right), along with estimates from WaTL and Only Target methods.
-[…]
+To further illustrate the effectiveness of WaTL, we visualize the cumulative distribution function of physical activity levels for one female participant and one male participant in Figure 2(b), along with estimates from WaTL and local Fréchet regression, using only the target data.
+The results indicate that WaTL provides a better fit to the true distribution, outperforming the estimate obtained using only the target data.
+We further evaluate the robustness of WaTL on a human mortality dataset in Appendix 8, where WaTL continues to demonstrate strong performance.
 7 Conclusion
 We introduce Wasserstein transfer learning and its adaptive variant for scenarios where the informative set is unknown, addressing the challenges posed by the lack of linear operations in the Wasserstein space.
-By leveraging the Wasserstein metric, the proposed algorithm accounts for the non-Euclidean structure of distributional outputs, ensuring compatibility with the geometry of the Wasserstein space.
+By leveraging the Wasserstein metric, the proposed algorithm accounts for the non-Euclidean structure of distributional outputs, ensuring compatibility with the intrinsic geometry of the Wasserstein space.
 Supported by rigorous theoretical guarantees, the framework demonstrates improved estimation performance compared to methods that rely solely on target data.
@@
 2015).
-Extending the methodology to higher dimensions presents an exciting future direction and would involve addressing both computational and theoretical challenges specific to high-dimensional Wasserstein spaces.
-Our theoretical analysis assumes that each probability distribution is fully observed, consistent with much of the existing literature on distributional data analysis (Petersen & Müller 2019, Chen et al.
-2023).
-However, this assumption could be unrealistic in practical applications, where only independent samples from the underlying distributions are available.
-This limitation can be addressed by replacing the unobservable distributions with empirical measures constructed from these samples (Zhou & Müller 2024).
-The derived convergence rates can be extended to incorporate an additional term reflecting the number of independent samples per distribution, providing a straightforward way to account for this practical consideration.
-[…]
-We further assess WaTL using the age-at-death distributions from 162 countries in 2015, compiled from the United Nations Databases [^2] https://data.un.org/ and the UN World Population Prospects 2022333https://population.un.org/wpp/Download.
-The dataset provides country-level age-specific death counts, which we convert into smooth age-at-death densities using local linear smoothing; see Figure 3 (a) for an illustration.
-For this analysis, we define the 24 developed countries as the target site and the remaining 138 developing countries as the source site.
-Given the limited sample size, we input the entire dataset into the model without performing cross-validation.
-Country-level age-at-death distributions are shaped by a range of factors, including economic conditions, healthcare systems, and social and environmental contexts.
-Among these, we select population density and gross domestic product (GDP) per capita as key predictive covariates.
-The results, shown in Figure 3 (b), demonstrate that the proposed WaTL method successfully transfers knowledge across these demographically diverse regions, effectively addressing the socioeconomic disparities between the source and target domains.
-Figure 3: (a) Age-at-death densities of developed and developing countries; (b) Root mean squared prediction risk (RMSPR) of WaTL and Only Target methods for human mortality data.
+Extending the methodology to higher dimensions is an exciting future direction that entails addressing both computational and theoretical challenges specific to high-dimensional Wasserstein spaces.
+Acknowledgments
+We thank the Area Chair and the reviewers for their constructive feedback. Doudou Zhou was supported by the NUS Start-Up Grant (A-0009985-00-00) and the MOE AcRF Tier 1 Grant (A-8003569-00-00).